OpenStudy (anonymous):

A rectangular barge 6 m long and 3 m wide floats in fresh water. How much will the barge sink when loaded with 600 kg of sand? Show all calculations leading to an answer.

OpenStudy (anonymous):

Let sink depth=d volume of displaced water ...V = 3x 6 x d V = 18 d .... is its weight..So, 18000d kg equals 600 kg.... so d= 0.033 metres @quick24

OpenStudy (anonymous):

@aryandecoolest hey are you still online

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

how did you change 18d to 18000 d?

OpenStudy (anonymous):

volume is in m^3 and density of water is 1000 kg/m^3

OpenStudy (anonymous):

you can do it other way round. The displaced water occupies about 600/1000=.6m³ .6m³/18m². Both gives you same answer :) hope it's clear

OpenStudy (anonymous):

ok so 600kg is the space the sand occupies in waters place and v=3x6xd because d=sink depth and we know 18d=v and v=600kg

OpenStudy (anonymous):

so therefor 600kg=18d

OpenStudy (anonymous):

you know buoyancy. Buoyancy,h=mg/pglw, l,w = length and width so h= 6000kg *9.8 m/s^2/1000kg/m^3*9.8m/s^2*6m*3m

OpenStudy (anonymous):

h=5880 kg *m/s^2/176400 kg/s^2=1/30m Hope it's quite clear.